# Dynamical control in a prethermalized molecular ultracold plasma: Local dissipation drives global relaxation

### v1

@@ III.2 Selective field ionization as a gauge of plasma electron binding energy as a function of time
The top frame of Figure 3 displays the contour of plasma SFI spectra formed by 4000 $44f(2)$ Rydberg gas volumes of systematically varying initial density obtained for a ramp delay of 2 $\mu$s. Note how Rydberg gases of high initial density avalanche fully, while those of lower density evolve to form a mixture of weakly bound NO+ ions and electrons together with a trace of Rydberg molecules with binding energies in the range of $n_{0}$, indicated by the dim signal in the bottom right of the SFI spectrum. An evident absence of electron signal forms a slight wedge at very low SFI potential in the upper left-hand corner of this contour. At the highest density, field ionization requires a minimum of a few V cm-1, which corresponds to Coulomb binding energy on the order of 300 GHz.
@@ III.6 Driven dissipation in the ultracold plasma as a function of mm-wave pulse width and delay
Figure 9: mm-wave driven dissipation of the ultracold plasma that evolves from a $42f(2)$ Rydberg gas, as driven by a mm-wave pulse tuned to the $42\ell(2)\rightarrow 43d(2)$ transition at 93.07 GHz, as a function of pulse width, measured going forward for a pulses that begin at $t_{\omega_{2}}$ (red) , and measured backwards from a pulse that ends at $t_{\rm G_{2}}$ (blue).
@@ IV.2 Selective field ionization and the initial dynamics of molecular ultracold plasma evolution
The SFI spectrum of a state-selected NO Rydberg gas measured with no ramp-field delay after $\omega_{2}$ excitation produces a broad Gaussian distribution of single-shot integrated SFI intensities, in which the binding energy spectrum largely consists of a simple pair of features reflecting the field-ionization threshold of the selected $44f(2)$ high Rydberg state with respect to the $N^{+}=0$ and 2 rotational states of the NO+ ion [108, 94].
By contrast, after a ramp delay of no more than a few $\mu$s, the SFI spectrum shows little, if any, signature of the initial Rydberg gas. As shown in Figure 3, electrons accounting for most of the signal in the 2 $\mu$s delayed SFI spectrum diabatically bind to $N^{+}=2$ rotational-state NO+ ions by a vertical energy of no more than 540 GHz. A light-blue region in the lower right of Figure 3 signifies a small residue of Rydberg molecules with the initially selected principal quantum number. The shift of this signal to a slightly higher appearance potential in the first 200 ns of evolution directly signals the effect of $\ell$-mixing electron collisions that drive residual Rydberg molecules to states of high electron orbital angular momentum. Hundreds of measurements such as these at each of 50 selected $\omega_{1}-\omega_{2}$ delays from 0 to 1000 ns yield wide distributions of integrated intensities, reflecting individually and in aggregate, a power law avalanche size distribution suggestive of scale invariance and a self-organized critical phase [94].
@@ V.5 Dimensional analysis
VI Conclusion
The molecular ultracold plasma that forms following the avalanche and quench of a state-selected Rydberg gas of nitric oxide exhibits an arrested relaxation that presents the signature of an enduring prethermal phase. These dynamics begin with double-resonant excitation, which prepares an initial ensemble composed of molecules in a single selected state of principal quantum number, $n_{0}$. Penning ionization, followed by an avalanche of electron-Rydberg collisions, forms a plasma of NO+ ions and weakly bound electrons, in which electron collisions drive a residual population of $n_{0}$ Rydberg molecules to a state of very $-\ell$. Predissociation further depletes this plasma of $-\ell$ molecules, purifying the $-\ell$ ensemble. For a wide range of initial conditions, this system enters a critical phase in which a density of NO+ and electrons balances a population of Rydberg molecules.

### v2

@@ III.2 Selective field ionization as a gauge of plasma electron binding energy as a function of time
The top frame of Figure 3 displays the contour of plasma SFI spectra formed by 4000 $49f(2)$ Rydberg gas volumes of systematically varying initial density obtained for a ramp delay of 6 $\mu$s. Note how Rydberg gases of high initial density avalanche fully, while those of lower density evolve to form a mixture of weakly bound NO+ ions and electrons together with a trace of Rydberg molecules with binding energies in the range of $n_{0}$, indicated by the dim signal in the bottom right of the SFI spectrum. An evident absence of electron signal forms a slight wedge at very low SFI potential in the upper left-hand corner of this contour. At the highest density, field ionization requires a minimum of a few V cm-1, corresponding to Coulomb binding energy on the order of 300 GHz.
@@ III.6 Driven dissipation in the ultracold plasma as a function of mm-wave pulse width and delay
Figure 9: mm-wave driven dissipation of the ultracold plasma that evolves from a $42f(2)$ Rydberg gas, as driven by a mm-wave pulse tuned to the $42\ell(2)\rightarrow 43d(2)$ transition at 93.07 GHz, as a function of pulse width, measured going forward for a pulses that begin at $t_{\omega_{2}}$ (blue) , and measured backwards from a pulse that ends at $t_{\rm G_{2}}$ (yellow).
@@ IV.2 Selective field ionization and the initial dynamics of molecular ultracold plasma evolution
The SFI spectrum of a state-selected NO Rydberg gas measured with no ramp-field delay after $\omega_{2}$ excitation produces a broad Gaussian distribution of single-shot integrated SFI intensities, in which the binding energy spectrum largely consists of a simple pair of features reflecting the field-ionization threshold of the selected $49f(2)$ high Rydberg state with respect to the $N^{+}=0$ and 2 rotational states of the NO+ ion [108, 94].
By contrast, after a ramp delay of no more than a few $\mu$s, the SFI spectrum shows little, if any, signature of the initial Rydberg gas. As shown in Figure 3, electrons accounting for most of the signal in the 6 $\mu$s delayed SFI spectrum diabatically bind to $N^{+}=2$ rotational-state NO+ ions by a vertical energy of no more than 540 GHz. A light-blue region in the lower right of Figure 3 signifies a small residue of Rydberg molecules with the initially selected principal quantum number. The shift of this signal to a slightly higher appearance potential in the first 200 ns of evolution directly signals the effect of $\ell$-mixing electron collisions that drive residual Rydberg molecules to states of high electron orbital angular momentum. Hundreds of measurements such as these at each of 50 selected $\omega_{1}-\omega_{2}$ delays from 0 to 1000 ns yield wide distributions of integrated intensities, reflecting individually and in aggregate, a power law avalanche size distribution suggestive of scale invariance and a self-organized critical phase [94].
@@ V.5 Dimensional analysis
VI Conclusions
The molecular ultracold plasma that forms following the avalanche and quench of a state-selected Rydberg gas of nitric oxide undergoes an arrested relaxation that exhibits the signature of an enduring prethermal phase. These dynamics begin with double-resonant excitation, which prepares an initial ensemble composed of molecules in a single selected state of principal quantum number, $n_{0}$. Penning ionization, followed by an avalanche of electron-Rydberg collisions, forms a plasma of NO+ ions and weakly bound electrons, in which electron collisions drive a residual population of $n_{0}$ Rydberg molecules to a state of very $-\ell$. Predissociation further depletes this plasma of $-\ell$ molecules, purifying the $-\ell$ ensemble. For a wide range of initial conditions, this system enters a critical phase in which a density of NO+ and electrons balances a population of Rydberg molecules.